Three charges were present inside the wire. Charge #1 has 3.2c, charge #2 has 4.1c, and charge #3 has -2.9c. Each charge has a distance of 8m to each other.


a. What is the total force between charges #1 and #2?

b. What is the total force between charges #2 and #3?

c. What is the total force and direction of the charges assuming that our charge #1 is the starting point? 


1
Expert's answer
2021-03-26T11:27:52-0400

Answer

Force between two charged particle

"F=\frac{kqq'}{r^2}"

a. Force between first and second charge

"F_{12}=\frac{9\times10^9\times3.2\times4.1}{(8)^2}=1.8\times10^9N"

b. Force between second and third charge

"F_{23}=\frac{9\times10^9\times2.9\times4.1}{(8)^2}=1.67\times10^9N"

Now

c. Force between first and third charge

"F_{13}=\frac{9\times10^9\times3.2\times2.9}{(16)^2}=0.33\times10^9N"

So net force on first charge

"F_{1}=F_{12}+F_{13}"

"=1.8\times10^9(-i) +0.33\times10^9(i)"

"=-1.47\times10^9(i)" N

In neagtive X-axis.